OpenStudy (anonymous):

The amount of money in an account with continuously compounded interest is given by the formula A= Pe^rt, where P is the principl, R is the annual interest rate, and t is time in years. Calculate to the nearest tenth of a year how long it take for an amount of a year how long it takes for an amount of money to double if interest is compounded continuously at 7.5% A. 8.2 years B. 9.2 years C. 0.9 years D. 0.6 years

OpenStudy (anonymous):

ima go with c but not 100%

OpenStudy (perl):

you can solve the equation 2P = P* e^(rt) , plug in r = .075

OpenStudy (anonymous):

2p= Pe^(0.5)(t) ?

OpenStudy (perl):

almost

OpenStudy (perl):

2P= Pe^(0.075*t)

OpenStudy (anonymous):

ohhh. sorry

OpenStudy (perl):

no problem :)

OpenStudy (perl):

now you can divide both sides by P, that will eliminate P

OpenStudy (anonymous):

why is P on both sides?

OpenStudy (perl):

the original equation was A = Pe ^(r * t)

OpenStudy (anonymous):

yeah but you wrote..."2P= Pe^(0.075*t)"

OpenStudy (perl):

right, because the question said, when does your principal double

OpenStudy (perl):

how long will it take for an initial amount P , to double. double P is 2P

OpenStudy (anonymous):

ohhh okay.

OpenStudy (perl):

you can also eliminate P by multiplying both sides of the equation by P

OpenStudy (anonymous):

2=Pe(0.075*t)

OpenStudy (perl):

should be 2 = e^(.075 * t )

OpenStudy (anonymous):

okay. got it.

OpenStudy (anonymous):

do i just plug in now?

OpenStudy (anonymous):

yeah thats correct

OpenStudy (wolf1728):

I'd say the first thing to do would be to convert the interest rate to a CONTINUOUSLY compounded rate. The formula for that is: http://www.1728.org/rate.htm Continuous rate = (e^annual rate) -1 Continuous rate = (2.718281828 ^ .075) -1 Continuous rate = 0.0778841509 Since we need to calculate the time for doubling, we'll say the initial amount is $100 and the ending amount is $200. The compounding formula solved for years is: http://www.1728.org/compint2.htm Years = {log(total) -log(Principal)} / log(1 + rate) Years = {log(200) -log(100)} / log(1+ 0.0778841509) Years = (2.3010299957 - 2) / 0.0325720861 Years = .3010299957 / 0.0325720861 Years = 9.2419624207 Looks like it is "b" don't you think?

OpenStudy (wolf1728):

There is a simpler way (but NOT as "dead on" exact) to calculate this. The APPROXIMATE method to calculate the "doubling time" of an interest rate is called the "Rule of 72". You get the number 72 and divide it by the interest rate. Doubling time in years = 72/annual rate Doubling time in years = 72/7.5 Doubling time in years = 9.6 We can easily see that "b" is the only answer that is even remotely close to 9.6. Of course if you require dead on accuracy, then the previous method is the one to use.

OpenStudy (wolf1728):

Upon further consideration, we must consider what the calendar tells us ... and it says: "Merry Christmas"

OpenStudy (anonymous):

I actually remember that now!

OpenStudy (wolf1728):

Glad I could help out. (Luckily Santa helped me as he dropped by my house.) :-)